# Zeroth-Order Policy Gradient for Reinforcement Learning from Human Feedback without Reward Inference

### v1

@@ 3 Preliminaries
Trajectory Reward: we assume the expected reward of each trajectory $\tau$ is a general function $r(\tau)$ which maps any trajectory to a value in $[0,H]$ (Zhang et al., 2024a). Without loss of generality, we scale the average per-step reward into $[0,1]$. Note that this model captures a broadened class of RL problems including classic MDPs, convex MDPs (Zhang et al., 2020; Zahavy et al., 2021) and maximum entropy exploration (Hazan et al., 2019). For any given policy $\pi$, we can formulate the initial value function $V_{1}^{\pi}(s)$ as the expected reward of trajectories starting from $s$ with policy $\pi$:
@@
where $\tau_{1}\succ\tau_{0}$ is the event that the human feedback prefers $\tau_{1}$ over $\tau_{0}.$ The human feedback $o,$ therefore, is a random sample from a Bernoulli distribution with ${\mathbb{P}}(o=1)={\mathbb{P}}(\tau_{1}\succ\tau_{0})$. The notion of link function comes from the dueling bandit literature to model preference with latent utility between arms, e.g. see Bengs et al. (2021, Section 3.2). This preference model has been used in dueling bandits (Bengs et al., 2021; Yue and Joachims, 2009; Kumagai, 2017; Ailon et al., 2014) as well as RLHF (Wang et al., 2023). One can see that one specific link function $\sigma$ will define a specific preference model, i.e., replacing $\sigma$ with a logistic function, we recover the Bradley-Terry model (Bradley and Terry, 1952), which is commonly used in RLHF for both practical (Christiano et al., 2017; Ouyang et al., 2022; Rafailov et al., 2023) and theoretical works (Du et al., 2024; Zhan et al., 2024a, b). The following assumption on the link function is standard in both dueling bandits (Bengs et al., 2021) and preference-based RL (Wang et al., 2023). One can easily verify that the Bradley-Terry model satisfies this assumption.
@@
The link function $\sigma(\cdot)$ in the preference model in equation 2 is bounded within $[0,1]$ and strictly monotonically increasing on $[-H,H]$ with $\sigma(0)=\frac{1}{2}$.
Problem Formulation: Our goal is to find policy network parameter ${\bm{\theta}}$ that maximizes the value function, i.e., $\max_{{\bm{\theta}}\in{\mathbb{R}}^{d}}V(\pi_{{\bm{\theta}}}).$
Notations: For a scalar $a$, we use $\mathrm{trim}[a|\Delta]$ to represent $\min\{\max\{a,\Delta\},1-\Delta\}$. For a vector ${\bm{v}}$, $\mathrm{trim}[{\bm{v}}|\Delta]$ represents the vector after applying the $\mathrm{trim}$ operator to each element respectively. Let ${\bm{e}}_{i}\in{\mathbb{R}}^{d}$ represent the unit vector with all zero elements but $1$ on the $i$-th coordinate.
@@ 4 Zeroth-Order Policy Gradient Algorithms for RLHF
In this section, we propose two RLHF algorithms without reward inference, motivated by the relationship between human feedback and zeroth-order gradient. We first present ZPG, which is a stochastic gradient descent algorithm and then ZBCPG, which is a stochastic block-coordinate descent algorithm.
@@ 4.1 ZPG: Zeroth-Order Policy Gradient from Human Feedback
8             estimate comparison probability $p_{t,n}=\mathrm{trim}\left[\left.\sum_{m=1}^{M}\frac{o_{n,m}}{M}\right|\Delta\right]$;
@@
Our first algorithm Zeroth-Order Policy Gradient from Human Feedback (ZPG) consists of the following five steps in each policy gradient iteration:
@@
The pseudo-code is presented in Alg. 1. As we mentioned earlier, our approach uses human feedback in a way different from both the classic reward inference in RLHF and DPO. The reward inference uses human preferences to recover the global reward function and DPO relates human preferences to the optimal policy. We view the human feedback as local information that points to the direction of a more preferred policy, i.e., the policy gradient direction. The algorithm has two key components: (i) a value function difference estimator from human preference and (ii) a policy gradient estimator from value function difference.
Policy Gradient Approximation: At each iteration of the algorithm, it first samples a $d$ dimensional vector ${\bm{v}}_{t}$ from a unit sphere and then perturbs the policy network parameter ${\bm{\theta}}_{t}$ along the direction of this sampled vector. Then, it uses the inner for loop to construct an estimation of the value function difference between the original policy and the perturbed policy, i.e., $V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}})$. We then plug it in the zeroth-order SGD algorithm proposed in (Ghadimi and Lan, 2013) to construct a zeroth-order approximate to the policy gradient, i.e.,
@@
$V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}})\approx\sum_{n=1}^{N}\sigma^{-1}(p_{t,n})/N.$ (3)
Querying humans multiple times ensures an accurate estimation of the reward gap between two trajectories. The reward gap of two trajectories is a random sample of the value function difference so we sample multiple trajectories to ensure the average trajectory reward gap converges to the value function difference. To ensure finite variance after applying $\sigma^{-1}$ function, we trim $p_{t,n}$ with a small constant which can be set to $\min\{\sigma(-H),1-\sigma(H)\}$ in this case.
@@ 4.2 ZBCPG: Zeroth-Order Block-Coordinate Policy Gradient from Human Feedback
9             estimate comparison probability $p_{t,n}=\mathrm{trim}\left[\left.\sum_{m=1}^{M}\frac{o_{n,m}}{M}\right|\Delta\right]$;
10
@@
13
Algorithm 2 Zeroth-Order Block Coordinate Policy Gradient from Human Feedback
In high-dimensional optimization problems, it is usually memory and operation inefficient to approximate the full gradient and update all the parameters in the policy network at the same iteration step (Malladi et al., 2023; Zhang et al., 2024b), which motivates parameter efficient fine-tuning (PEFT). The stochastic (block) coordinate descent approach naturally arises because of its ease of implementation, low memory requirements, and adaptability to distributed settings (Nesterov, 2012; Lu and Xiao, 2015). The same advantage also applies to RLHF when the number of parameters in the policy network is too large. Therefore, we propose a block coordinate version of ZPG, called Zeroth-Order Block Coordinate Policy Gradient (ZBCPG), which is summarized in Alg. 2. The key difference between ZBCPG and ZPG is the choice of the perturbation direction.
@@
$\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\approx\mathbb{E}_{{\bm{i}}_{t},\bm{\lambda}_{t}}\left[\frac{d}{K\mu}\left(V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})\right){\bm{v}}_{t}\right].$
The block-coordinate approach allows us to (i) perturb a subset of parameters at each iteration, e.g., a specific layer of the policy network for fine-turning, and (ii) have a parallel implementation where we have multiple gradient estimators $\hat{{\bm{g}}}_{t}$ when updating the policy. We will later show that both algorithms have similar provable convergence guarantees but the analysis of ZBCPG is more challenging due to the perturbation mechanism.
@@ 5.1 Assumptions
In order to infer the local reward difference from human preference probability through link function $\sigma(\cdot)$, we impose the following assumption which is satisfied by the Bradley-Terry model. A slightly weaker assumption is also adopted by Wang et al. (2023) and justified as a minimal requirement to learn the optimal policy. We use $\Delta=\min\{\sigma(-H),1-\sigma(H)\}$ as the trim constant.
@@
Since a trajectory reward is bounded in $[0,H]$, $V(\pi_{{\bm{\theta}}^{*}})<\infty$, where ${\bm{\theta}}^{*}$ is the global optimal solution. For simplicity, we assume the same $L$ satisfies both assumptions above.
@@ 5.2 Convergence Rate and Sample Complexity
In this section, we present the theoretical guarantees for both ZPG and ZBCPG under all three assumptions above. We aim to learn an ${\epsilon}$-stationary policy, i.e., a policy $\pi_{{\bm{\theta}}}$ with $\|\nabla_{{\bm{\theta}}}V(\pi_{\bm{\theta}})\|_{2}^{2}\leq{\epsilon}$, and study the convergence rate and sample complexity of the value function gradient.
@@
Choose the perturbation distance and learning rate to be
@@
If $M=\Omega(D^{2})$ and we randomly pick ${\bm{\theta}}_{R}$ uniformly from the trajectory $\{{\bm{\theta}}_{0},{\bm{\theta}}_{1},\cdots,{\bm{\theta}}_{T-1}\}$, then the convergence rate of ZPG satisfies:
@@
Choose the perturbation distance and learning rate to be
@@
If $M=\Omega(D^{2})$ and we randomly pick ${\bm{\theta}}_{R}$ uniformly from the trajectory $\{{\bm{\theta}}_{0},{\bm{\theta}}_{1},\cdots,{\bm{\theta}}_{T-1}\}$, then the convergence rate of ZBCPG satisfies:
@@
$\underbrace{\frac{Hd}{T}}_{\text{Zeroth-Order GD}}+\underbrace{\frac{d^{2}\sqrt{\log M}}{\sqrt{M}}}_{\text{Preference Estimation}}+\underbrace{\frac{Hd\sqrt{d}}{\sqrt{N}}.}_{\text{Value Function Approximation}}$
The second represents the error that occurs when using multiple human preferences $[o_{n,1},\cdots,o_{n,M}]$ to approximate the population-level human preference probability for given two trajectories, i.e., ${\mathbb{P}}(\tau_{n,1}\succ\tau_{n,0})$. This error will further result in a bias term after being plugged into the inverse link function $\sigma^{-1}(\cdot)$ to construct an estimation of the value function difference. The third term comes from the variance of using multiple trajectory rewards to approximate the value function of a policy. The first term representing the error resulting from zeroth-order stochastic gradient descent or blocked coordinate descent, which matches the state-of-the-art analysis result ${\mathcal{O}}(d/T)$ for non-convex smooth function optimization (Nesterov and Spokoiny, 2017). Even though the final convergence rates are the same and we both use constant learning rates, how we choose the perturbation distance to obtain the rate differs from (Ghadimi and Lan, 2013). Specifically, they chose a small perturbation distance with $\mu^{2}={\mathcal{O}}(d/T)$ to make sure the zeroth order approximation error is of lower order. However, this choice will not work for us, because our gradient estimate is biased due to the non-linear nature of the link function in preference estimation. If we choose the perturbation distance $\mu$ to be too small, the preference estimation error will be amplified by $d/\mu$ due to the formula of zeroth-order approximation $\hat{{\bm{g}}}_{t}$. This phenomenon adds complication to our theoretical analysis. Our method is to use a moderate perturbation distance $\mu$. Moreover, this moderate perturbation distance also balances the preference estimation and the value function approximation errors.
Based on the two theorems above, we have the following corollary that characterizes the sample complexity of our proposed algorithms.

### v2

@@ 2 Preliminaries
In this section, we introduce the notations and preliminaries of our paper. We first introduce the RL problem with general trajectory reward. Then, we describe the policy parameterization, and the human preference feedback mechanism, as well as our optimization problem formulation.
@@
Trajectory Reward: we assume the expected reward of each trajectory $\tau$ is a general function $r(\tau)$ which maps any trajectory to a value in $[0,H]$ (Zhang et al., 2024a). Without loss of generality, we scale the average per-step reward into $[0,1]$, and the return of the trajectory does not necessarily need to be the sum of per-step rewards. For any given policy $\pi$, we can formulate the initial value function $V_{1}^{\pi}(s)$ as the expected reward of trajectories starting from $s$ with policy $\pi$:
@@
where $\tau_{1}\succ\tau_{0}$ is the event that the human feedback prefers $\tau_{1}$ over $\tau_{0}.$ The human feedback $o,$ therefore, is a random sample from a Bernoulli distribution with ${\mathbb{P}}(o=1)={\mathbb{P}}(\tau_{1}\succ\tau_{0})$. The notion of link function comes from the dueling bandit literature to model preference with latent utility between arms, e.g. see Bengs et al. (2021, Section 3.2). This preference model has been used in dueling bandits (Bengs et al., 2021; Yue and Joachims, 2009; Kumagai, 2017; Ailon et al., 2014) as well as RLHF (Wang et al., 2023). One can see that one specific link function $\sigma$ will define a specific preference model (Azari et al., 2012), i.e., replacing $\sigma(\cdot)$ with a logistic function, we recover the Bradley-Terry model (Bradley and Terry, 1952), which is commonly used in RLHF for both practical (Christiano et al., 2017; Ouyang et al., 2022; Rafailov et al., 2023) and theoretical works (Du et al., 2024; Zhan et al., 2024a, b). On the other hand, let $\sigma(\cdot)$ be the cumulative distribution function (CDF) of standard normal distribution, we obtain the Probit model (Thurstone, 1927) which has wide application in the study of social choice theory, economy, and psychology (Train, 2009; Greene, 2000). A detailed discussion on other preference models is provided in Section 1.2. The following assumption on the link function is standard in both dueling bandits (Bengs et al., 2021) and preference-based RL (Wang et al., 2023). One can easily verify that both the Bradley-Terry and Probit models satisfy this assumption.
@@
The link function $\sigma(\cdot)$ in the preference model in equation 2 is bounded within $[0,1]$ and strictly monotonically increasing on $[-H,H]$ with $\sigma(0)=1/2$.
Problem and Notations: Our goal is to find parameter ${\bm{\theta}}$ that maximizes the value function, i.e.,
$\max_{{\bm{\theta}}\in{\mathbb{R}}^{d}}V(\pi_{{\bm{\theta}}}).$
For a scalar $a$, we use $\mathrm{trim}[a|\Delta]$ to represent $\min\{\max\{a,\Delta\},1-\Delta\}$. For a vector ${\bm{v}}$, $\mathrm{trim}[{\bm{v}}|\Delta]$ represents the vector after applying the $\mathrm{trim}$ operator to each element respectively. Let ${\bm{e}}_{i}\in{\mathbb{R}}^{d}$ represent the unit vector with all zero elements but $1$ on the $i$-th coordinate.
@@ 3 Zeroth-Order Policy Gradient Algorithms for RLHF
In this section, we propose two RLHF algorithms without reward inference, motivated by the relation between preference and zeroth-order gradient. We first present ZPG, a stochastic gradient descent algorithm, and then ZBCPG, a stochastic block-coordinate descent algorithm.
@@ 3.1 ZPG: Zeroth-Order Policy Gradient from Human Feedback
estimate preference probability:
[…]
@@
Our first algorithm ZPG consists of the following five steps in each policy gradient iteration:
@@
The pseudo-code is presented in Alg. 1. As we mentioned earlier, our approach uses human feedback in a way different from both the classic reward inference in RLHF and DPO. The reward inference uses human preferences to recover the global reward function and DPO relates the human preference generation mechanism to the optimal policy. We view the human feedback as local information that points to the direction of a more preferred policy, i.e., the policy gradient direction. Some online RLHF algorithms such as online DPO also exploit similar local estimation viewpoints, i.e., using new trajectories of current policy to locally improve the estimation of DPO loss and then proceed. However, we want to emphasize that the relation between the DPO loss and the optimal policy is still global, and it is limited to deterministic MDPs, which shows the novelty of our approach. The algorithm we propose has two key components: (i) a value function difference estimator from human preference and (ii) a policy gradient estimator from value function difference.
Policy Gradient Approximation: At each iteration of the algorithm, it first samples a $d$ dimensional vector ${\bm{v}}_{t}$ from a unit sphere and then perturbs the policy network parameter ${\bm{\theta}}_{t}$ along the direction of this sampled vector. Then, it uses the inner for loop to construct an estimation of the value function difference between the original policy and the perturbed policy, i.e., $V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}})$. We then plug it into the zeroth-order stochastic gradient descent (SGD) algorithm proposed in (Ghadimi and Lan, 2013) to construct a zeroth-order approximation to the policy gradient, i.e.,
@@
$V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}})\approx\frac{1}{N}\sum_{n=1}^{N}\sigma^{-1}(p_{t,n}).$ (3)
Querying humans multiple times ensures an accurate estimation of the reward gap between two trajectories. The reward gap of two trajectories is a random sample of the value function difference, so we sample multiple trajectories to ensure the average trajectory reward gap converges to the value function difference. To ensure finite variance after applying $\sigma^{-1}(\cdot)$ function, we trim $p_{t,n}$ with a small constant which can be set to $\min\{\sigma(-H),1-\sigma(H)\}$ in this case.
Remarks on Local and Global Reward Estimation. The main philosophical difference between our proposed algorithms and classic RLHF paradigms is the use of local reward information around the neighborhood of the current policy. This is completely different from RLHF with reward inference since reward models intend to approximate the global reward landscape of the MDP. Some recently developed online RLHF algorithms, such as online DPO (Guo et al., 2024) and XPO (Xie et al., 2024) also employ certain local reward information to estimate the DPO loss in each iteration. Specifically, the algorithms will explore the trajectories in the neighborhood of the current policy to construct a local estimation of the DPO loss. However, on the other hand, the relation between the DPO loss and the best policy is still global in deterministic MDPs studied in both papers. Our paper’s approach is purely local and does not rely on such global relations, which is incorrect in stochastic MDPs. It estimates the local policy gradient directly from human feedback, and proceeds gradient descent for policy updates. The use of local information in our approach is more general and may apply to a wider range of RL problems.
@@ 3.2 ZBCPG: Zeroth-Order Block-Coordinate Policy Gradient from Human Feedback
9             estimate preference probability:
$p_{t,n}=\mathrm{trim}\left[\left.\sum_{m=1}^{M}\frac{o_{n,m}}{M}\right|\Delta\right];$
@@
12
Algorithm 2 Zeroth-Order Block-Coordinate Policy Gradient from Human Feedback
In high-dimensional optimization problems, it is usually memory restricted or operation inefficient to approximate the full gradient and update all the parameters in the policy network at the same iteration step (Malladi et al., 2023; Zhang et al., 2024b), which motivates parameter efficient fine-tuning (PEFT). The stochastic (block) coordinate descent approach naturally arises because of its ease of implementation, low memory requirements, and adaptability to distributed settings (Nesterov, 2012; Lu and Xiao, 2015). The same advantage also applies to RLHF when the number of parameters in the policy network is too large. Therefore, we propose ZBCPG, a block coordinate version of ZPG, which is summarized in Alg. 2. The key difference between ZBCPG and ZPG is the choice of the perturbation direction, where we use Rademacher noise instead of the normal perturbation in ZPG.
@@
$\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\approx\mathbb{E}_{{\bm{i}}_{t},\bm{\lambda}_{t}}\left[\frac{d}{K\mu}\left(V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}})\right){\bm{v}}_{t}\right].$
The block-coordinate approach allows us to (i) perturb a subset of parameters at each iteration, e.g., a specific layer of the policy network for fine-tuning, and (ii) have a parallel implementation where we have multiple gradient estimators $\hat{{\bm{g}}}_{t}$ when updating the policy. We will later show that both algorithms have similar provable convergence guarantees but the analysis of ZBCPG is more challenging due to the perturbation mechanism.
Remarks on KL Regularization in RLHF: the KL-regularization is commonly used in RLHF for both empirical and theoretical studies, e.g., PPO (Ouyang et al., 2022), DPO (Rafailov et al., 2024a, b; Guo et al., 2024; Dong et al., 2024), and more recent algorithms (Xie et al., 2024; Xiong et al., 2024). On the positive side, the KL-regularization improves the training stability in RLHF tasks. For example, in LLM applications, the model pre-training procedure will provide powerful capabilities such as reasoning, and the regularization ensures that the updated policy is unlikely to deviate from the pre-training model and lose these capabilities. It also ensures the global relation between the optimal policy and the reward function in deterministic MDPs for LLM studies and motivates easier objectives such as the DPO loss. On the other hand, in general RL problems, if the regularization is too large, it also limits the RL agent from the possibility of learning the best policy, which may be very different from the reference policy of the pre-trained model. It is also worth noting that KL-regularized MDPs may be easier to explore compared to their unregularized counterparts, especially for MDPs with sparse rewards. This is because the best policies for regularized MDPs are likely to be stochastic policies which exhibit certain exploration power due to stochasticity. In contrast, the optimal policy for unregularized counterparts is deterministic, which does not explore. It seems there is no golden rule on whether KL-regularization is indeed needed in RLHF applications beyond LLMs. The answer is likely to depend on specific applications. Our algorithm can also be applied to KL-regularized MDPs after a slight modification in gradient estimation. In this case, the zeroth-order gradient of the policy will include the sum of two components, the value function difference estimation, and the KL regularization difference, i.e.,
$\left(V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-\beta\mathsf{KL}(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}}\|\pi_{{\bm{\theta}}_{0}})\right)-\left(V(\pi_{{\bm{\theta}}_{t}})-\beta\mathsf{KL}(\pi_{{\bm{\theta}}_{t}}\|\pi_{{\bm{\theta}}_{0}})\right),$
since the objective is now the KL-regularized value function. Here, $\pi_{{\bm{\theta}}_{0}}$ is the reference policy, and $\mathsf{KL}(\pi_{\bm{\theta}}\|\pi_{{\bm{\theta}}_{0}})$ is the KL divergence between a policy parameterized by ${\bm{\theta}}$ and the reference policy, i.e.,
$\mathsf{KL}(\pi_{\bm{\theta}}\|\pi_{{\bm{\theta}}_{0}})=\mathbb{E}_{s_{h},a_{h}\sim\pi_{\bm{\theta}}}\left[\sum_{h=1}^{H}\log\frac{\pi_{\bm{\theta}}(a_{h}|s_{h})}{\pi_{{\bm{\theta}}_{0}}(a_{h}|s_{h})}\right].$
Remark that human evaluators will only provide preference based on the trajectory returns, not the KL regularization. Therefore, the value function difference component can be estimated the same way as the current algorithm, i.e., line $9$ in Alg. 1. Then, the rest is to estimate the KL regularization difference between the perturbed policy $\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}}$ and the current policy $\pi_{{\bm{\theta}}_{t}}$, concerning the reference policy $\pi_{{\bm{\theta}}_{0}}$, i.e.,
$\mathsf{KL}(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}}\|\pi_{{\bm{\theta}}_{0}})-$ $\mathsf{KL}(\pi_{{\bm{\theta}}_{t}}\|\pi_{{\bm{\theta}}_{0}})$
$=$ $\mathbb{E}_{s_{h},a_{h}\sim\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}}}\left[\sum_{h=1}^{H}\log\frac{\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}}(a_{h}|s_{h})}{\pi_{{\bm{\theta}}_{0}}(a_{h}|s_{h})}\right]-\mathbb{E}_{s_{h},a_{h}\sim\pi_{{\bm{\theta}}_{t}}}\left[\sum_{h=1}^{H}\log\frac{\pi_{{\bm{\theta}}_{t}}(a_{h}|s_{h})}{\pi_{{\bm{\theta}}_{0}}(a_{h}|s_{h})}\right].$
This can be achieved by estimating the KL regularization term of both policies, which is conducted by evaluating the action probability of the reference policy $\pi_{{\bm{\theta}}_{0}}$ on the trajectories generated from both policies. With both components, we can construct the zeroth-order gradient for the regularized MDP and conduct gradient descent similar to ZPG and ZBCPG.
@@ 4.1 Assumptions
To infer the local reward difference from human preference probability through link function $\sigma(\cdot)$, we impose the following assumption which is satisfied by the Bradley-Terry model. A slightly weaker assumption is also adopted by Wang et al. (2023) and justified as a minimal requirement to learn the optimal policy. We use $\Delta=\min\{\sigma(-H),1-\sigma(H)\}$ as the trim constant.
@@
Since a trajectory reward is bounded in $[0,H]$, $V(\pi_{{\bm{\theta}}^{*}})<\infty$, where ${\bm{\theta}}^{*}$ is the global optimal solution. For simplicity, we assume $L$ is the constant upper bound for both assumptions.
@@ 4.2 Convergence Rate and Sample Complexity
In this section, we present the theoretical guarantees for both ZPG and ZBCPG under all three assumptions mentioned in previous sections. We aim to learn an ${\epsilon}$-stationary policy $\pi_{{\bm{\theta}}}$ with $\|\nabla_{{\bm{\theta}}}V(\pi_{\bm{\theta}})\|_{2}^{2}\leq{\epsilon}$, and study the convergence rate and sample complexity.
@@
Choose the perturbation distance $\mu$ and the learning rate $\alpha$ to be chosen as follows:
@@
If $M=\Omega(H^{2})$ and we randomly pick ${\bm{\theta}}_{R}$ uniformly from the trajectory $\{{\bm{\theta}}_{0},{\bm{\theta}}_{1},\cdots,{\bm{\theta}}_{T-1}\}$, then the convergence rate of ZPG satisfies:
@@
Choose the perturbation distance $\mu$ and the learning rate $\alpha$ to be chosen as follows:
@@
If $M=\Omega(H^{2})$ and we randomly pick ${\bm{\theta}}_{R}$ uniformly from the trajectory $\{{\bm{\theta}}_{0},{\bm{\theta}}_{1},\cdots,{\bm{\theta}}_{T-1}\}$, then the convergence rate of ZBCPG satisfies:
@@
$\underbrace{\frac{Hd}{T}}_{\text{Zeroth-Order Gradient Descent}}+\underbrace{\frac{d^{2}\sqrt{\log M}}{\sqrt{M}}}_{\text{Preference Estimation}}+\underbrace{\frac{Hd\sqrt{d}}{\sqrt{N}}.}_{\text{Value Function Approximation}}$
The second represents the error that occurs when using multiple human preferences $[o_{n,1},\cdots,o_{n,M}]$ to approximate the population-level human preference probability for given two trajectories, i.e., ${\mathbb{P}}(\tau_{n,1}\succ\tau_{n,0})$. This error will further result in a bias term after being plugged into the inverse link function $\sigma^{-1}(\cdot)$ to construct an estimation of the value function difference. The third term comes from the variance of using multiple trajectory rewards to approximate the value function of a policy. The first term represents the error resulting from zeroth-order stochastic gradient descent or blocked coordinate descent, which matches the state-of-the-art analysis result ${\mathcal{O}}(d/T)$ for non-convex smooth function optimization (Nesterov and Spokoiny, 2017). Even though the final convergence rates are the same and we both use constant learning rates, how we choose the perturbation distance to obtain the rate differs from (Ghadimi and Lan, 2013). Specifically, they chose a small perturbation distance with $\mu^{2}={\mathcal{O}}(d/T)$ to make sure the zeroth order approximation error is of lower order. However, this choice will not work for us, because our gradient estimate is biased due to the non-linear nature of the link function in preference estimation. If we choose the perturbation distance $\mu$ to be too small, the preference estimation error will be amplified by $d/\mu$ due to the formula of zeroth-order approximation $\hat{{\bm{g}}}_{t}$. This phenomenon adds complication to our theoretical analysis. Our method is to use a moderate perturbation distance $\mu$. Moreover, this moderate perturbation distance also balances the preference estimation and the value function approximation errors.
Based on the theorems, we have the following corollary that characterizes the sample complexity.
@@
Remarks on Local and Global Convergence. Our results only establish the local convergence of both proposed algorithms. In general, considering the generality of stochastic MDPs and policy parameterization studied in our paper, it is extremely difficult to establish global convergence results without further assumptions beyond the ones that have already been made in the paper, such as smoothness. We conjure that our proposed algorithms will have global convergence under additional assumptions on the value function, such as convexity or the Polyak-Lojasiewicz (PL) condition (Karimi et al., 2016). Again, these assumptions are often questionable in practice. Some contemporary works, such as Xie et al. (2024), established global convergence in the deterministic KL-constrained MDPs with additional coverability assumptions, a much easier and narrowed model where the optimal policy can be explicitly expressed. Considering the difficulty and generality of our setting, we believe local convergence is the best result that can be obtained, and a meaningful result showing the effectiveness of the proposed algorithm to a wider range of RL problems beyond LLMs. It remains an open question of how to design algorithms that achieve global optimality under minimal and practical assumptions.